OpenStudy (anonymous):

how many liters of peroxide should be added to 12 liters of an 8% peroxide solution so that the remaining solution contains 16% peroxide?

OpenStudy (tkhunny):

Equate Something! P is the amount of peroxide to add. Equate Peroxide: 12*0.08 + P = (12+P)*0.16 Equate NonPeroxide: 12*0.92 = (12+P)*0.84